24 October 2009 LBC Tech Night 2

We took a lot of data last night and much still needs to be analyzed, but we learned some things right away.

1) Red Z22 zero-point is just about right, Blue Z22 zero-point needs a -170 nm correction.

2) The technical chip focus corrections were not working as they should.

As a result:

  • Andrea B changed the FPIA active optics code and uploaded a new version to lbccontrol. The Trackers PC needs to be rebooted to pick up this change. At the start of the night I rebooted the CMU (there was an error which necessitated a reboot) and restarted LBC.

  • We need to change Blue Z22 zero point from 2.3. With this current value, there is about 170nm too much spherical on the "collimated" pupil. Need to determine by what amount to change Zero Point (in lbcfpia.cfg).


Plan for tonight:

  • Test active optics from tech chip. Go to NGC110 and run for 30 min.

19:08 Wayne cleared active optics, added +1000nm Z11 to left. He also gave -2000nm Z7 to both SX and DX. FPIA tried to send another -2000 Z7, hit -Y limit. Z7 <-> Y. So cntl c'ed and restarted. OK now.

19:30 Slew to NGC110. Focus/collimate and observe with 15-min OB. Monitor "Active Optics P03" in lbc.log and Z4 sent to both LEFT/RIGHT PSF GUI. In case we need later to check messages logs: PCS on 302; PSF LEFT on 303; PSF RIGHT on 303; PMC LEFT 301; PMC RIGHT 304.

Blue guide star out of focus and Red guide star astigmatic during exposure. And on science image, red has 6 pix fwhm and blue, noticeably out of focus and fwhm ~10pix.

No active optics corrections were sent, it seems: no output when grepping for Active Optics P03 in log...

in lbc.log we see:
2009/10/24 02:58:04.669000 N R TRACKERS IA       "lbciarun.exe" program completed
2009/10/24 02:58:04.685000 W R TRACKERS IA       error reading "c:\lbcfpia\src\lbciaCoeffs_red.dat" file
2009/10/24 02:58:06.362913 W R TRACKERS          IA System error [src/trackers/trackers.c:1728]
2009/10/24 02:58:06.362970 N R TRACKERS          Active Optics corrections calculation error

2009/10/24 02:58:09.809000 N B TRACKERS IA       "lbciarun.exe" program completed
2009/10/24 02:58:09.809000 W B TRACKERS IA       error reading "c:\lbcfpia\src\lbciaCoeffs_blue.dat" file
2009/10/24 02:58:07.985771 W B TRACKERS          IA System error [src/trackers/trackers.c:1728]
2009/10/24 02:58:07.985833 N B TRACKERS          Active Optics corrections calculation error

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -2570   -193    221      7     56    525    100    y  1.35  1.07                      /newdata/lbcb.20091024.031208.fits  V-BESSEL
 -1435    122   -606    148   -127    147    100    y  1.44  0.61                      /newdata/lbcb.20091024.031353.fits  V-BESSEL
  -633     31    184     11    110    110     67    y  1.30  0.28                      /newdata/lbcb.20091024.031601.fits  V-BESSEL
  -686      4    274     37    -23     90      9    y  1.20  0.30                      /newdata/lbcb.20091024.031805.fits  V-BESSEL
  -420    155   -222    -21      4     76     11    y  1.25  0.21                      /newdata/lbcb.20091024.032012.fits  V-BESSEL
  1399      0      0      0      0   -736      0    y  0.00  0.00                     dofpia_sph3_adjustment             


 -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1189    -30    334    343     80    -92      7    y  0.89  0.51                      /newdata/lbcr.20091024.031207.fits  r-SLOAN 
  -483      0   -229    -30     45     63      2    y  0.98  0.22                      /newdata/lbcr.20091024.031347.fits  r-SLOAN 
  -409    132    102    -90   -102      6     -6    y  0.98  0.20                      /newdata/lbcr.20091024.031556.fits  r-SLOAN 
  -255   -153    -16    111     -3    -33    -13    y  0.79  0.15                      /newdata/lbcr.20091024.031800.fits  r-SLOAN 
  -294     83   -355    -36     59    -23     -4    y  0.85  0.18                      /newdata/lbcr.20091024.032008.fits  r-SLOAN 
   707      0      0      0      0   -372      0    y  0.00  0.00                     dofpia_sph3_adjustment             

20:22 Running 30 min OB on NGC110... Guide stars for blue/red are much tighter and ~1" or less. We know (see above) that the corrections are not being sent, so this can be considered as a baseline.

-1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -309    301   -334    231     10     16   -100    y  1.06  0.22                      /newdata/lbcb.20091024.035612.fits  V-BESSEL
  1272      0      0      0      0   -669      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -255   -117    220    301     90    -19    -51    y  0.63  0.20                      /newdata/lbcr.20091024.035612.fits  r-SLOAN 
   562      0      0      0      0   -296      0    y  0.00  0.00                     dofpia_sph3_adjustment             

04:11 Changed lbcfpia.cfg file to drop BlueZ22Zero from 2.3 to 1.5. With 2.3, "converged' pupils have Z22, they show a ring around the outer perimeter and ring around the inner hole. Running dofpia with this new limit:

-1140 0 0 0 0 600 0 y 0.00 0.00 dofpia_sph3_adjustment
  1. -491 492 -139 105 150 -100 y 1.08 0.33 /newdata/lbcb.20091024.041410.fits V-BESSEL
  2. 196 -157 90 39 -35 -100 y 1.03 0.17 /newdata/lbcb.20091024.041550.fits V-BESSEL 1254 0 0 0 0 -660 0 y 0.00 0.00 dofpia_sph3_adjustment

04:21 dofpia had converged. I reinstated the 100nm threshold for Z22. I had increased this to 101 so that we would not keep sending 100 nm of Z22, but now that we changed the Zero Point, this may not happen. Running dofpia again:
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   328    199   -422     86    -13   -263   -100    y  1.03  0.27                      /newdata/lbcb.20091024.042141.fits  V-BESSEL
   731   -152   -232    -74      7   -172    -61    y  1.02  0.33                      /newdata/lbcb.20091024.042323.fits  V-BESSEL
  -174     24   -392      1     27    -67    -26    y  1.12  0.16                      /newdata/lbcb.20091024.042533.fits  V-BESSEL
  1316      0      0      0      0   -692      0    y  0.00  0.00                     dofpia_sph3_adjustment             

04:33 The resulting 'converged' pupil after this change of blue Z22 zeropoint looks flatter. Saved as lbtdu16:/home/LBTO/BlueZ22Zero_2.3vs1.5.jpg

04:35 There is a chip or bite at about 40 degrees counter-clockwise from the spider. Taking a 4-chip extrafocal pupil image to see if the orientation changes. lbcb.20091024.043306.fits. Chip size doesn't seem to be dependent on location in field. We are taking an extra/intra focal pair now. 044228 (B extra) and 044318 (B intra). Another pair 045727(B extra) and 045816(B intra)

Wayne examined the LBC -Blue L1 and SX mirror but saw nothing that might be causing this chip. This seems to be a new feature - don't know from when, though.

05:56 focus/collimate on NGC110 as before. Take extra/intra to check if Z11 is still present in blue. FPIA estimates seeing about 1.2" Blue and 0.8" Red... or from 1-1.2 Blue and 0.6-0.8 Red. Blue 055702 and 055752. Z11 is still there.

06:19 We experimented with the Z11 BlueFinalZ11Zero, changing it from -300 to 0.0. Going to see if we still have spherical in the ex/in focal pairs. So to summarize here we have Z22 zeropoint = 1.5 (vs 2.3) and Z11 finalzero = 0 (vs -300)

extra: 062409 (B) and 062405 (R) --- these show Z11 but less than previous set (055702...) intra: 062459 (B) and 062616 (R)

06:30 Changed Z11 BlueFinalZ11Zero to increase it still futher, from 0 to +150. Running dofpia, and now taking ex/in focal pupils:

extra: 063617 (B) and 063612 (R) intra: 063706 (B) ...

Based on examining the sequence (see /home/LBTO/Z11FinalZero_sequence_150_0_-350.jpg)... The 6 panels, left to right, top to bottom are

extra and intra focal pairs for Final Z11 = 150, 0 and the original -300. All are with Z22ZeroPoint=1.5.

06:50 We changed the Z11 BlueFinalZ11Zero to +300 and are running dofpia in prep for taking an extra/intra focal sequence.

extra: 065827 (B) and 065821 (R) intra: 065917 (B) and 065935 (R) (Right main mirror failed to focus).

07:05 Not too much difference between +300 and +150, but +300 is ok - leave at that as BlueFinalZ11Zero.

07:05 Now go back to original values of both BlueZ22Zero (1.5 -> 2.3) and BlueFinalZ11Zero (300 -> -300). Running dofpia and then going to take extra/intra focal pupils...

extra: 071539 (B) and 071532 (R) intra: 071630 (B) and 071616 (R)

07:25 Left BlueFinalZ11Zero at -300, the original value, and changed BlueZ22Zero to 1.5.

extra: 074358 (B) intra: 074449 (B)

This too appears to have spherical... I made a ds9 image with, from left to right, top to bottom:

All Blue (code for image Z11andZ22progressions.jpg)

063617 063706 062409 062459
055702 055752 065827 065917
071539 071630 074358 074449

08:00 Slew to new field: NGC1023, which is high. Andrea Baruffolo is online - and checking tech chip code. First try --- guide stars were astigmatic... loads of astig. We are running dofpia again to remove this.
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   259    980    675     20    127    258     -2    y  0.86  0.41                      /newdata/lbcb.20091024.081817.fits  V-BESSEL
   -31   -433     97     40      8    -52     -9    y  1.00  0.16                      /newdata/lbcb.20091024.081944.fits  V-BESSEL
   382    659    126    221    156     45      8    y  0.80  0.29                      /newdata/lbcb.20091024.082141.fits  V-BESSEL
   421    348    154    -73    182    -80      8    y  0.82  0.24                      /newdata/lbcb.20091024.082348.fits  V-BESSEL
  -243   -219   -188     10   -157     94    -17    y  0.91  0.18                      /newdata/lbcb.20091024.082548.fits  V-BESSEL
  1172      0      0      0      0   -617      0    y  0.00  0.00                     dofpia_sph3_adjustment             

08:28 Starting OB (blue only due to earlier HP problem on Red --- Wayne since Panicked the mirror and it is recovered now, but since I focused only on Blue, am running this OB with Blue only).

08:48 Wayne panicked DX and brought it back up, so I ran dofpia again on both sides. Both sides took a while to converge as astigmatism was bouncing around for both --- cut/pasting the logs so you can see if this might be an incidence of telescope vibration, although at the outset it doesn't seem the red/blue are in sync.

  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  4050   1247    147  -2701  -1428    247     28    y  0.00  2.14                      /newdata/lbcr.20091024.083736.fits  r-SLOAN 
  -811   -946      8    452    393   -238    -19    y  0.63  0.53                      /newdata/lbcr.20091024.083949.fits  r-SLOAN 
  -567     59    -74     46     57    156     -9    y  0.59  0.26                      /newdata/lbcr.20091024.084300.fits  r-SLOAN 
  -108     48    133     27      0    117      6    y  0.59  0.12                      /newdata/lbcr.20091024.084517.fits  r-SLOAN 
   536      0      0      0      0   -282      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   179    108    222    -19     68    162      1    y  0.68  0.17                      /newdata/lbcb.20091024.083728.fits  V-BESSEL
   105   -781   -800    113    -11   -271     42    y  1.03  0.38                      /newdata/lbcb.20091024.083954.fits  V-BESSEL
  -728   1066    929    -43    229    440     10    y  0.84  0.59                      /newdata/lbcb.20091024.084305.fits  V-BESSEL
   912   -174   -324    142     85   -332     -8    y  0.82  0.44                      /newdata/lbcb.20091024.084522.fits  V-BESSEL
  1112      0      0      0      0   -585      0    y  0.00  0.00                     dofpia_sph3_adjustment             

09:00 I changed the Z22 Zero Point for Blue back to 2.3. NOTE: 1.5 is the better value as far as Z22, but to use it we need to determine new filter focus offsets ---- the previous few frames on NGC1023 (taken after the focus run up to 084522) confirm this. Reran dofpia and am now redoing the OB on NGC1023 to resume testing of the active optics corrections from tech chips.

090014..090323... 150 sec exposures. Guide star fwhm = 1.5 (B) and 0.8 (R) then later 1.1 and 0.9... While Red guide star fwhm is stable around 0.9"-1", the blue is large at times... 1-1.8".

09:33 OB completed and running dofpia again on field...
 -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -875     41    136     45      6   -130    -40    y  0.52  0.37                      /newdata/lbcr.20091024.093304.fits  r-SLOAN 
  -167   -151     65    -79   -102    -20    -13    y  0.53  0.13                      /newdata/lbcr.20091024.093442.fits  r-SLOAN 
   495      0      0      0      0   -261      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1052   -253    -72     42     93    249     50    y  1.00  0.46                      /newdata/lbcb.20091024.093308.fits  V-BESSEL
  -378    222   -119     25      6     96     17    y  0.98  0.19                      /newdata/lbcb.20091024.093446.fits  V-BESSEL
  1219      0      0      0      0   -641      0    y  0.00  0.00                     dofpia_sph3_adjustment             

09:45 Slewing to open cluster NGC 1749... We didn't have much Active Optics luck with NGC1023 last night either. Silent Red controller error meant we lost the Red test (3sec) image... Tried to rerun an OB, but it hung at the red controller error. Turning off other systems... but it won't --- really seems to be hung. Finally came up with error box and allowed me to turn off other systems... From lbc.log
2009/10/24 09:50:36.318604 E R CAMERA            reinitializing CCD Controller failed
2009/10/24 09:50:36.318725 E R CAMERA            timeout on CCD Controller [src/camera/camera.c:1534]
2009/10/24 09:50:36.318865 E R                   camera system encountered a problem
2009/10/24 09:50:36.319232 N R                   preset completed
2009/10/24 09:50:36.319259 E R                   >>> preset/expose/save interrupted at preset time
2009/10/24 09:50:37.518678 W R CAMERA            connection timeout [src/camera/camera.c:969]
2009/10/24 09:50:37.518820 W R CAMERA            connection timeout [src/camera/camera.c:969]
2009/10/24 09:50:37.578862 N R                   completed 1 OB image/s
2009/10/24 09:50:37.578892 E R                   OB execution error: none [src/lbc/actions.c:1603]
2009/10/24 09:50:37.578975 N -                   freeing FITS headers
2009/10/24 09:50:37.579432 N -                   UI update with message >(RED Channel) OB execution error: (CAMERA) CCD Controller error!< at 1256377837.579211
2009/10/24 09:50:37.579696 N -                   UI update at 1256377837.579553
2009/10/24 09:50:37.579801 N -                   executing PLAY command completed
2009/10/24 09:50:38.718777 W R TRACKERS          connection timeout [src/trackers/trackers.c:1267]
2009/10/24 09:50:38.868949 N R CAMERA            closing connection for a while
2009/10/24 09:50:38.869023 N R TRACKERS          closing connection for a while


2009/10/24 09:54:19.065642 E R CAMERA            RPC Client/Host communication failed retval:5 [src/camera/camera.c:820]
2009/10/24 09:54:19.065797 E R                   camera system encountered a problem

2009/10/24 09:59:14.443000 E R CAMERA   CCDCTRL  "CCD Controller fiber link error or incomplete data" (code:22) controller error [ccdctrl.c:2417]

2009/10/24 09:59:18.771000 E R CAMERA   CCDCTRL  "SPC not found - reboot please" (code:18) controller error [ccdctrl.c:2378]
2009/10/24 09:59:18.771000 E R CAMERA   CCDCTRL  hardware problems [ccdctrl.c:1373]

10:20 Finished running dofpia on NGC1749 field - and now running an OB. Guide stars look good: 0.8" both B and R.

Blue image (102023) has FWHM ~ 3.34 (0.77") g-SLOAN and Red 3.2 (0.7") r-SLOAN.

11:00 OB ended. Andrea thinks he found the source of error but cannot fix it at this time. Ran dofpia again after OB ended:
  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   807   -215    -10    345     32   -208     14    y  0.65  0.39                      /newdata/lbcr.20091024.105359.fits  r-SLOAN 
    21     25    216   -160     30   -154      1    y  0.52  0.16                      /newdata/lbcr.20091024.105545.fits  r-SLOAN 
   490      0      0      0      0   -257      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   264   -159      5    314    -89   -157     51    y  0.91  0.23                      /newdata/lbcb.20091024.105403.fits  V-BESSEL
  -371   -246     27    -17     48    123     -6    y  0.86  0.20                      /newdata/lbcb.20091024.105549.fits  V-BESSEL
  1140      0      0      0      0   -600      0    y  0.00  0.00                     dofpia_sph3_adjustment             

11:03 Going to investigate filter focus offsets for Blue when Z22 Zero Point is changed.

BlueZ22Zero = 2.3; RB_rV_superfoc: 110350... (0.0046 with fwhm=3.40) and then RB_ig_superfoc: 110758... (0.052 with FWHM=3.0)

11:15 now BlueZ22Zero = 1.5, running dofpia....
 -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
    84    -47   -132    248   -156   -169      9    y  0.53  0.19                      /newdata/lbcr.20091024.111620.fits  r-SLOAN 
  -246      2    151    -60    -69    -55      0    y  0.52  0.14                      /newdata/lbcr.20091024.111802.fits  r-SLOAN 
  -123     36   -121   -107    -26     18     11    y  0.62  0.11                      /newdata/lbcr.20091024.112031.fits  r-SLOAN 
  -225    -87    -78     68    118     36      1    y  0.56  0.14                      /newdata/lbcr.20091024.112233.fits  r-SLOAN 
   517      0      0      0      0   -272      0    y  0.00  0.00                     dofpia_sph3_adjustment     
        
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -269   -105     90     22     79     21   -100    y  0.93  0.14                      /newdata/lbcb.20091024.111624.fits  V-BESSEL
   272     37   -232     38    -38   -177   -100    y  0.92  0.19                      /newdata/lbcb.20091024.111806.fits  V-BESSEL
   439     40    259      0    -19   -132    -39    y  0.92  0.22                      /newdata/lbcb.20091024.112035.fits  V-BESSEL
   -43     -9   -214     -2    -16    -43    -27    y  0.91  0.11                      /newdata/lbcb.20091024.112237.fits  V-BESSEL
  1174      0      0      0      0   -618      0    y  0.00  0.00                     dofpia_sph3_adjustment

RB_rV_superfoc: 112505... (-0.055 with FWHM=3.86 this is 0.06mm below canonical focus offset) and then RB_ig_superfoc: (0.03 ... new lower limit, with FWHM=3.57; so this is a difference of 0.05mm below). RB_rV_superfocZ22 and RB_ig_superfocZ22 starting focus values 0.06 lower.

and RB_YU while analyzing: 113339... 113611... (0.273 again at lower limit... fwhm=3.61 --- nominal was 0.353 )

11:37 Changed lbcfpia.cfg again to set BlueZ22Zero = 2.3, running dofpia...

RB_YU: 114543 (0.309 with FWHM=3.96 -- so this is 0.036~0.04 above result with new Z22 zeropoint));

RB_RUs 114922...(0.394 with FWHM = 3.53... nominal was 0.435).

RB_zB: (0.023 with fwhm=4.04... nominal was 0.053).

RB_Ir:115827...(0.0522 with FWHM=3.13 --- nominal was 0.086)

12:00 Change BlueZ22Zero =1.5 and run superfoc for Us, B and r.

RB_RUs_Z22: ...120835...121030.121108 (0.408 with FWHM = 3.16... this is close to the value just obtained with Z22 Zero=2.3).

RB_zB_Z22: 121231...121502: (0.0174 with FWHM = 3.24... again, this is close to the value just obtained with Z22 Zero = 2.3).

RB_Ir_Z22: 122106...122337 (not a good fit --- Z22 apparent).

12:22 Run dofpia again for Zero=1.5 -- blue converged on first iteration. This time RB_rV...
  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   767     89    511   -362     86   -178     47    y  1.18  0.40                      /newdata/lbcr.20091024.122641.fits  r-SLOAN 
  -357    -44   -265    400    -94     98     10    y  0.81  0.26                      /newdata/lbcr.20091024.122932.fits  r-SLOAN 
  -372   -230   -204     64     -5    244    -14    y  0.74  0.25                      /newdata/lbcr.20091024.123144.fits  r-SLOAN 
   634      0      0      0      0   -333      0    y  0.00  0.00                     dofpia_sph3_adjustment     
        
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   -63   -161   -226    104     -8     25    -40    y  1.01  0.13                      /newdata/lbcb.20091024.122506.fits  V-BESSEL
   269   -149    245     52    -36     -3     -4    y  1.17  0.16                      /newdata/lbcb.20091024.122936.fits  V-BESSEL
    32    242    123     46     -1    -50    -29    y  1.11  0.12                      /newdata/lbcb.20091024.123147.fits  V-BESSEL
  1307      0      0      0      0   -688      0    y  0.00  0.00                     dofpia_sph3_adjustment     

RB_rV_superfoc for Z22: 123357...(looks like it could be around -0.06, or consistent with earlier measurement but it is not a good fit).

At end of night, changed BlueZ22Zero in lbcfpia.cfg to original 2.3. Ran dofpia and took another RB_rV superfoc:

RB_rV_superfoc: 124502... (-0.05...)

set of biases (note with dome open and at twilight - looks dark on webcam still 05:51 MST)

13:00 flats at g and i PA=0 and PA=180 (5sec).

13:06 Closing. I turned off LBC from User Interface: Other systems and HouseKeeping.


Summary of the Night:

A short summary:

1 Andrew & I edited the lbcfpia.cfg file BlueZ22Zero to change this Blue Z22 ZeroPoint from 2.3 to 1.5. It looks better, but this change has consequences to Z11 and to the filter focus offsets.

2 Andrea Baruffolo connected just before midnight and we tested the active optics computations from the tech chips. We worked for a few hours, but he found the source of the problem but could not address it during the night --- would there be a chance of using LBC in the next few nights?

3 Finally, I did try to determine filter focus offsets with the two different Z22 Zero Points at Blue. Initially it looked like the change in Zero Point from 2.3->1.5 resulted in a drop in focus offset by 0.06mm, but later sequences were not consistent with this and suggested even the same focus offset. I need to analyze these data more.

-- OlgaKuhn - 24 Oct 2009
I Attachment Action Size Date Who Comment
BlueZ22Zero_2.3vs1.5.jpgjpg BlueZ22Zero_2.3vs1.5.jpg manage 26 K 24 Oct 2009 - 05:41 OlgaKuhn Comparison of "converged" (last dofpia) pupil for Blue Z22 ZeroPoint=2.3 (left) and 1.5 (right)
Blue_SX_obscuration.jpgjpg Blue_SX_obscuration.jpg manage 23 K 24 Oct 2009 - 05:44 OlgaKuhn Obscuration at about 8 or 4 o'clock on LBC Blue lens or SX mirror?
Z11FinalZero_sequence_150_0_-350.jpgjpg Z11FinalZero_sequence_150_0_-350.jpg manage 61 K 24 Oct 2009 - 06:57 OlgaKuhn Changing BlueFinalZ11Zero from +150, 0, -300 (extra/intra focal pairs from left to right, top to bottom)
Z11andZ22progressions.jpgjpg Z11andZ22progressions.jpg manage 97 K 24 Oct 2009 - 08:06 OlgaKuhn See table which lists image numbers - all lbcb.20091024.??????.fits
Topic revision: r2 - 24 Oct 2009, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback